OpenStudy (anonymous):

I GIVE MEDALS!!!!! What is the equation of the line that passes through the point (3, 2) and has a slope of 4? A.) y = 4x - 1 B.) y = 4x - 5 C.) y = 4x - 10 D.) y = 4x + 1

OpenStudy (anonymous):

Just substitute the point in the eqution which eqution satisfies that is the answer....

OpenStudy (anonymous):

did you get it....@turtleluver

OpenStudy (anonymous):

im thinking either a or c...but im not sure

OpenStudy (anonymous):

that is nothing but 1) 2=4(3)-1 not satisfied..... and so on....

OpenStudy (anonymous):

Yes you are exactly correct Option C is correct...

OpenStudy (anonymous):

ty

OpenStudy (the_fizicx99):

The formula is: y-y1=m(x-x1) y-2=4(x-3) can you solve now?

OpenStudy (anonymous):

Option C satisfies the point...

OpenStudy (anonymous):

That is the 2nd approch....@tHe_FiZiCx99 we can also do inthat way....